What is the average cost for linear search on equality on a primary key field? Assume 100,000 records, 10 records/block, 5 ms access time, 100 MB/s transfer rate, and 4 KB block size. Assume sequential access.
Answer #1

So, from the data we can say the total number of block= 100000/10= 10000

block size of 1 block= 4KB

block size of 10000 blocks=40000KB=40MB

Transfer Rate=100MB/s

Total time for transferring block will be= 40/100 sec= 4ms

Access time=5ms

Retrieve a single record that satisfies the corresponding equality condition

so, Cost = (hi + 1) * (tT + tS )=(1+1)*(5+4)=18